OpenStudy (anonymous):

A country's population in 1990 was 123 million.In 2002 it was 128 million.Estimate the population in 2013 using the exponential growth formula.(P=Ae^kt).Round to the nearest millionth..

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (anonymous):

@science0229

OpenStudy (anonymous):

P=Number of Population.. A=Number Of Population At Time..Which Is 0? K=Positive Constant.. E=The Natural Logarithm Base?

OpenStudy (anonymous):

Thats What I Have In My Notes.

OpenStudy (science0229):

How about t?

OpenStudy (anonymous):

t=0

OpenStudy (science0229):

t can't be zero, and neither can be A.

OpenStudy (science0229):

This is what I think. I think t is the year, and A and k are positive constants, which we have to solve for.

OpenStudy (anonymous):

Sorry For The Misunderstanding..A=Number Of Population A Time..Time,AKA T=0

OpenStudy (anonymous):

At The Time*

OpenStudy (science0229):

What do you mean?

OpenStudy (anonymous):

Do It Your Way,Im Very Unsuccessful I These Types Of Problems..

OpenStudy (science0229):

ok

OpenStudy (science0229):

So, you're given 2 positive variables and 2 data points.

OpenStudy (science0229):

Wait, sorry. 2 equations.

OpenStudy (anonymous):

Yes.

OpenStudy (science0229):

You're given 2 data points to solve for 2 variables, 'A' and 'k'. First equation is \[123=Ae ^{1990k}\] (I'll skip writing million for 123) Second equation is \[128=Ae ^{2002k}\]. If you divide second equation by first equation, you'll get\[\frac{ 128 }{ 123 }=e ^{12k}\].

OpenStudy (science0229):

By the definition of ln, \[12k=\ln \frac{ 128 }{ 123 }\]

OpenStudy (anonymous):

Alright.. So With e^12k,You Plug In e and k?

OpenStudy (science0229):

Divide both sides by 12, and you'll get \[k=\frac{ 1 }{ 12 }e ^{12}\]. That's k.

OpenStudy (anonymous):

Okay.

OpenStudy (science0229):

Now, you can plug that in for either of the equation, and find A using a calculator.

OpenStudy (science0229):

After you find A and k, you get a complete function that tells you the population in millions for a certain year, t.

OpenStudy (science0229):

So you plug in 2013 and round to the nearest millions.

OpenStudy (anonymous):

My Calculator Doesn't Support Fractions With Exponents..

OpenStudy (science0229):

Then you can try going to one of the online calculators.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I Got 134.232323?

OpenStudy (anonymous):

I Used This Calculator http://www.wolframalpha.com/

OpenStudy (science0229):

Just a second

OpenStudy (science0229):

Wait!!!

OpenStudy (science0229):

I'm really sorry.

OpenStudy (science0229):

k is supposed to be \[\frac{ 1 }{ 12 }\ln \frac{ 128 }{ 123 }\].

OpenStudy (science0229):

But that's basically how you do this problem.

OpenStudy (science0229):

Sorry, again.

OpenStudy (anonymous):

So After I Round,I Got 133? Correct?

OpenStudy (anonymous):

@science0229

OpenStudy (science0229):

if you got A and k correct, and you made no mistakes, that should be correct.

OpenStudy (science0229):

I'm not going to tell you the answer, but if you didn't make any calculation mistakes, that's the right answer.

OpenStudy (anonymous):

It Was Wrong..But Thanks Anyways,It Really Helped When You Explained It.Thanks :)

OpenStudy (science0229):

Welcome :)